OpenStudy (katielong):

Hey, i'm a little stuck on rearranging equations. I'll post the question below..

OpenStudy (katielong):

\[y=\frac{ 3(x-5) }{ 6-2x }\] make x the subject

OpenStudy (phi):

as a first step, multiply both sides by (6-2x) what do you get ?

OpenStudy (katielong):

y(6-2x)=3(x-5)

OpenStudy (phi):

The idea is you want the x's to all be "up top" (even if on the wrong side)

OpenStudy (phi):

next step, distribute the y on the left side and the 3 on the right side what do you get ?

OpenStudy (katielong):

by distribute, do you mean expand the brackets??

OpenStudy (phi):

The idea is to get the x's out of the parens. yes, distribute means expand the brackets.

OpenStudy (katielong):

if so, 6y-2yx=3x-15 ?

OpenStudy (phi):

next add -2yx to both sides (this gets the x's on the same side) add +15 to both sides what do you get ?

OpenStudy (phi):

oops, +2xy not -2xy

OpenStudy (katielong):

6y+15=3x+2yx

OpenStudy (phi):

any idea what to do next ?

OpenStudy (katielong):

ermmmmm... just gotta get that 'y' away so either divide by y or put in brackets or something??

OpenStudy (phi):

how about factor out the x so you get (stuff) * x then divide both sides by (stuff) to get x by itself?

OpenStudy (katielong):

so.. 6y+15=x(3+2y) and finally \[\frac{ 6y+15 }{ 3+2y}=x\]

OpenStudy (phi):

what you have is good, but it might be written differently often people would write the bottom as 2y+3 (putting the variable first) but good work!

OpenStudy (katielong):

ahh i see, yeh that's that order thing where theres x" then ?x then ? etc... yeh thanks alot :)

OpenStudy (katielong):

would \[\sqrt{\frac{ 5 }{ x+a }} \] be \[\frac{ 5 }{ Y ^{2} }-A\]

OpenStudy (phi):

yes on a multiple choice you might see \[ \frac{ 5- a y^2}{y^2} \] or \[ -\frac{ a y^2-5 }{y^2} \] if they changed a to a y^2/y^2 (to get a common denominator) and then added the two fractions. in the last form, you can "factor out" a -1 to that (5-ay^2) is written as -(ay^2-5) (this makes the ay^2 positive) of course \[ \frac{ -(a y^2-5) }{y^2} = -\frac{ a y^2-5 }{y^2} \]

OpenStudy (phi):

next time make it a new post, so you can get a medal.

OpenStudy (katielong):

haha okayy thanks, well i have a couple more questions, just writing them now :)